OpenStudy (anonymous):

a crew consists of person A person B and person C the crew receives a check for 8050 for a job they just finished if person B makes 300% of what person A makes, and person C makes 250% of what person B makes how much does each person make??

OpenStudy (anonymous):

B=3A C=2.5B=7.5A A+3A+7.5A=8050

OpenStudy (anonymous):

solve for A

OpenStudy (anonymous):

A=700 B=3A=700*3=2100 C=2.5*B= 2100* 2.5= 5250